the position of a particle is given by the expression x = 4cos (3πt+πi), where x is in meters and t is in seconds. Calculate the frequency, the period of the motion and the amplitude of motions.


1
Expert's answer
2022-03-20T18:47:39-0400

"\omega=3\pi~\frac{rad}s,"

"T=\frac{2\pi}{\omega}=0.67~s,"

"A=4~m."